1. Determine the accumulated value of P2000.00 in 5 years if it is invested at 11%

compounded quarterly. (Please show solution with correct answer)


1
Expert's answer
2022-04-16T05:34:15-0400

A = P(1 + r/n)nt

P: the principal, amount invested

A: the new balance

t: the time, years

r: the rate, (in decimal form)

n: the number of times it is compounded.

Compounded quarterlyn = 4:

A = 2000.00(1 + 0.11/4)(4)(5) = 2000(1.0275)(20) = $3440.86